I would appreciate it if someone can find the highlighted boxes, and explain what does the required return mean?
Clorox Dividend History PVs -year 1, FVe eyear 5, ne number of years-1, solve i Caterpillar Inc Clorox 2012 20 20 $1.962 $2.48 S2 Caterpillar Stock Valuation Stock values D 1 / (Required Return-Growth) Do Di Required return Growth Stock valuation Stock Price $3.08 18.0% $136.62 Clorox stock valuation Stock value D,/(Required Return-Growth) Do $3.14 D1 Required return Growth Stock valuation Stock Price 8.5% $128.11
Answer #1
Caterpillar
D1=Dividend of year1
D0=dividend in year 0
g= growth rate
D1=D0*(1+g)
R=Required rate of return
Price (P0)=D1/(R-g)
D0=dividend in year 0 $3.08
R= Required return=18%           0.18
Price=$136.62=(3.08*(1+g)/(0.18-g)
136.62*(0.18-g)=3.08*(1+g)
24.5916-136.62*g=3.08+3.08g 24.5916
139.7*g=24.5916-3.08 21.5116
g= growth rate=21.5116/139.7= 0.153984
Growth =15.4%
D1=3.08*(1+0.154)= $3.55
D0 $3.08
D1 $3.55
Required Return 18%
Growth 15.4%
Stock valuation=3.55/(0.18-0.154) $136.62
Clorox
D1=Dividend of year1
D0=dividend in year 0
g= growth rate
D1=D0*(1+g)
R=Required rate of return
Price (P0)=D1/(R-g)
D0=dividend in year 0 $3.14
R= Required return=8.5%         0.085
Price=$128.11=(3.14*(1+g)/(0.085-g)
128.11*(0.085-g)=3.14*g+3.14
10.88935-128.11*g=3.14*g+3.14 10.88935 7.74935
131.25*g=10.88935-3.14=7.74935 131.25
g= growth rate=7.74935/131.25= 0.059043
Growth =5.9%
D1=3.14*(1+0.059)= $3.33
D0 $3.14
D1 $3.33
Required Return 8.5%
Growth 5.9%
Stock valuation=3.33/(0.085-0.059) $128.11